# M3RS: Multi-robot, Multi-objective, and Multi-mode Routing and Scheduling

### v2

@@ Abstract
The quality of task execution can significantly impact a multi-robot mission. While higher quality is desirable, it may not always be feasible due to mission constraints. Existing multi-robot task allocation literature generally overlooks quality of service as a decision variable. Addressing this gap, we introduce the multi-robot, multi-objective, and multi-mode routing and scheduling (M3RS) problem, designed for time-bound, multi-robot, multi-objective missions. In M3RS, each task offers multiple execution modes, each with different resource requirements, execution time, and quality. M3RS optimizes task sequences and execution modes for each agent. The need for M3RS comes from multi-robot applications in which a trade-off between multiple criteria can be achieved by varying the task level quality of service through task execution modes. Such ability is particularly useful for service robot applications. We use M3RS for the application of multi-robot disinfection in healthcare environments and other public locations. The objectives considered for disinfection application are disinfection quality and number of tasks completed. A mixed-integer linear programming (MIP) model is proposed for M3RS. Further, a clustering-based column generation (CCG) algorithm is proposed to handle larger problem instances. Through synthetic, simulated, and hardware case studies, we demonstrate the advantages of M3RS, showing it provides flexibility and strong performance across multiple metrics. Our CCG algorithm generates solutions 2.5$\times$ faster than a baseline MIP optimizer, maintaining competitive performance. The videos for the experiments are available on the project website [^1] https://sites.google.com/view/g-robot/m3rs/.
Index Terms:
multi-robot system, autonomous systems, task allocation, scheduling, service robots, disinfection robots.
@@ I Introduction and Related Works
Maintaining hygiene in healthcare environments is crucial to prevent catastrophic issues like hospital-acquired infections. Disinfection robots are useful technologies that provide an additional layer of protection by automating the disinfection of target surfaces. A team of these robots can improve both spatial and temporal efficiency in maintaining hygiene standards in hospitals [1]. To coordinate a fleet of robots for a disinfection mission, two key decisions must be made: i) which task will be assigned to which robot, and ii) what level of disinfection quality should be applied to each task. While higher disinfection quality is desirable, it also requires more time and resources, which may not always be feasible. The assignment of tasks to robots has been studied in the literature [2], but there is a lack of research addressing the decision regarding the quality of service of a task. In disinfection applications, considering the quality of tasks is important as it impacts overall hygiene standards. Additionally, some surfaces, such as those in patient rooms and ICUs, require stringent quality standards. In this work, we tackle the problem of deciding which tasks to implement and at what mode/quality.
For a team of robots, Multi-robot Task Allocation (MRTA) algorithms [3, 2] allocate tasks to various robots working on a common mission while optimizing a joint mission objective. A key metric to evaluate these allocations is the quality of service (QoS) for the mission. The definition of QoS depends on the mission requirements and target application, such as the quality of communication between the robot and the base station [4, 5], the tardiness of addressing tasks [6], the average waiting time for a customer [7], or productivity and fatigue [8]. Often, there is a trade-off between QoS and other mission objectives, which could be conflicting. In such cases, multiple solutions exist that prioritize these objectives [4, 5, 8].
Figure 1: M3RS is suitable for routing and scheduling in time-bound scenarios where multiple criteria of interest depend on the mode in which the agents execute the tasks. This visualization depicts two cases. Here, mode 1 stands for the highest quality and long task execution times, whereas mode 3 is for the lowest quality and fastest task execution times. The left side of the image shows a case where agents choose to maximize the number of tasks completed by performing tasks in lower-quality modes. On the right, agents prioritize quality over quantity. As a result, tasks 2 and 5 are missed. An example of such a scenario is a multi-robot system performing services such as disinfection and cleaning.
MRTA problems are $\mathcal{NP}$-hard, making larger instances challenging to solve. With extra variables, concerning the quality of the service, the problem becomes more complex. Techniques such as heuristics and metaheuristics are faster but often sacrifice optimality [9]. Learning-based methods require expensive training and massive datasets, and may still struggle to generalize [10]. Column generation is a decomposition approach that refines the solution iteratively by solving smaller sub-problems. This allows faster convergence to higher-quality solutions as it directly exploits the problem’s structure and efficiently handles large-scale instances. We propose a clustering-based column generation algorithm, where clustering further helps in reducing the search space to find good quality solutions.
To the best of our knowledge, this is the first work to address task-level service quality using multiple execution modes for quality-dependent multi-robot missions. We introduce a multi-robot, multi-objective, and multi-mode routing and scheduling (M3RS) problem. The contributions of this work are:
@@
    We present M3RS, a novel problem to alter task-level QoS to achieve trade-offs in different mission objectives. The proposed problem is well-suited for applications where tasks are spread across a physical space, each task has service requirements and resource requirements, and the overall mission has multiple criteria of interest. Each robot can address the task in multiple execution modes, where each mode corresponds to different task quality, resource consumption, and service times. A visual depiction of the application of M3RS is given in Fig. 1.
@@
    We present the usage of M3RS for a multi-robot disinfection application. A thorough study of different mission scenarios is conducted on a synthetic dataset. Further, we test M3RS in simulated and hardware case studies.
@@
    We use a mixed-integer linear programming formulation for the proposed problem. To handle the large decision space, we propose clustering-based column generation (CCG) for solving M3RS. CCG generates solutions 2.5$\times$ faster than the baseline while maintaining competitive performance.
@@
II Literature Review: QoS in Task Assignemnt
QoS has also received attention in routing and production scheduling literature. Exposito et al. [11] considered the timeliness of vehicles as a QoS indicator for vehicle routing. Similarly, QoS has been studied for home healthcare applications, where it involves routing nurses to different patients for healthcare-related services, depending on the patient receiving care from a nurse with the desired experience [12, 13]. In supply chain and production planning, QoS has been considered as the existence of minimum stock [14]. Most of these existing works modify the task allocation to achieve different trade-offs between QoS and other objectives. However, there are no works that focus on the deliberate degradation or improvement of QoS offered to an individual task by a robot as a means to achieve an overall trade-off for various mission objectives.
Task execution modes are different operational settings for servicing tasks. These modes depend on various constraints in the environment or aim to achieve desired performance on key metrics like quality and resource consumption. Lahijanian et al. [15] explored using an energy-saving mode where resource-intensive autonomy features are disabled to conserve energy in resource-constrained environments. The usage of task execution modes in MRTA has been limited. Task execution modes were used to decide if humans and robots should collaborate to maximize time efficiency [16, 17]. In project scheduling, task execution modes are used to minimize project makespan while adhering to resource constraints [18, 19, 20]. In vehicle routing, mode selection involves choosing delivery locations with specific constraints, impacting customer delivery [21]. The hybrid electric vehicle routing problem in [22] used mode selection for switching driving modes, affecting emissions and transportation costs. Using task execution modes to select QoS provided to the task is still unexplored. This decision affects the quality of individual tasks and the overall mission. For example, in a disinfection mission, cleaning a surface with the highest available dosage ensures the highest quality level. However, higher dosages require longer service times, which could potentially result in skipping other tasks.
@@ III Multi-Robot Multi-Objective Multi-Mode Routing and Scheduling
In this section, first, the general optimization model for M3RS problem is presented. Then we present a column generation (CG) scheme to solve the optimization model with faster computation times. Our novel contribution is the use of clustering heuristics, which helps solve the problem more quickly and find solutions with better objective values. Finally, the formulation is applied to the multi-robot disinfection application.
@@
Notation Description
$\mathcal{A}$ Set of agents.
$\mathcal{T}$ Set of tasks.
$\mathcal{M}_{i}$ Set of modes for task $i$, where $\mathcal{M}_{i}=\{1,2...M\}$.
$R$ Number of objective criteria.
$f_{r}$ $r^{th}$ objective criterion.
$p_{i,m}$ The quality indicator of doing task $i$ in mode $m$.
$p^{max}$ The highest possible quality of disinfection
$\sigma_{i,m}$ Time to complete task $i$ in mode $m$.
$q_{i\scriptsize{\xrightarrow{}}j}$ Energy consumed for travelling from task $i$ to task $j$.
$q_{idle}^{k}$ Energy consumption during idling by agent $k$.
$q_{i,m}$ Energy consumption of task $i$ in mode $m$.
$\text{q}_{\text{travel}}$ Avg. current consumption in traveling at a constant speed.
$\text{q}_{\text{tc}}$ Avg. current consumption while completing tasks.
$\text{q}_{\text{idling}}$ Avg. current consumption while idling.
$[a_{i},b_{i}]$ Start and end time for task $i$.
$Q$ Total capacity of an agent.
$T_{H}$ Time available to complete all tasks.
$t_{i,j}$ Travel time from task $i$ to task $j$.
$\lambda_{r}$ The scalar preference for $r^{th}$ objective function.
$s^{k}_{i}$ A decision variable for arrival time of agent $k$ to task $i$.
$x^{k}_{i{\xrightarrow{}}j}$ A binary decision variable for if agent $k$ goes from task $i$ to task $j$.
$y^{k}_{i,m}$ A binary decision variable for task $i$ done in mode $m$ by agent $k$.
$\eta$ The load factor hat is used to simulate different disinfection missions.
@@ III-A Optimization Model
A solution for the M3RS problem determines routes, schedules, and task execution modes for a set of agents $\mathcal{A}$, and a set of tasks $\mathcal{T}$ distributed across physical space. Refer to Table I for notation. All the agents are identical, and only one agent can service each task. We use $i$ and $j$ to denote different tasks in $\mathcal{T}$ and $k$ to denote $k^{th}$ agent in the rest of the notation. The time window for a task $i$ specifies a starting time, $a_{i}$, and finishing time, $b_{i}$, within which the task must be serviced. The entire multi-robot operation must be done in a fixed mission time of $T_{H}$ hours. Further, each task $i$ has an associated mode set $\mathcal{M}_{i}=\{1,2...M\}$, representing available task service modes. Each task can have distinct modes of service available, so each task has a distinct mode set. An agent services the task in one of the modes available in the associated mode set. The selected mode $m\in\mathcal{M}_{i}$ for task $i$ has a corresponding service time ($\sigma_{i,m}$) and energy requirement ($q_{i,m}$) from the agent. Performing a task in a mode with higher service time and resource consumption would lead to lower availability of agents for other tasks. The overall scheduling and allocation mission is optimized for multiple objectives.
The energy capacity for each agent is given by $Q$ Ah (Ampere-hours). The energy consumption for each agent is divided into performing tasks, traveling, and idling [23]. A fixed-rate battery model is used to calculate the energy consumption, where energy consumed is proportional to idling time, task execution time, and traveling time respectively [23]. Different aspects of energy expenditure can be calculated using:
$\begin{split}q_{i\scriptsize{\xrightarrow{}}j}=\text{q}_{\text{travel}}\cdot t_{i,j},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\small(i.e. traveling energy)}\\ q_{i,m}=\text{q}_{\text{tc}}\cdot\sigma_{i,m},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\small(i.e. servicing energy)}\\ q_{idle}^{k}=\text{q}_{\text{idle}}\cdot(M_{k}-W_{k}).\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\small(i.e. idling energy)}\end{split}$ (1)
Here, $\text{q}_{\text{travel}}$ is the current consumption for an agent traveling at a constant speed, $\text{q}_{\text{tc}}$ is the current consumption while performing a task, and $\text{q}_{\text{idle}}$ is the current consumption for idling. $M_{k}$ is the total time spent by agent $k$ on the mission, and $W_{k}$ is the total time spent by agent $k$ traveling and executing tasks. The net energy consumed by agent $k$ in a mission is given by the LHS of Eq. (9). The relevant notation for the formulation is summarised in Table I.
@@
i) the binary decision variable $x^{k}_{i{\xrightarrow{}}j}\in\{0,1\}$ for assignment and transition of agent $k$ from task $i$ to $j$;
ii) the binary decision variable $y^{k}_{i,m}\in\{0,1\}$ for allocation of mode $m$ for task $i$ when executed by agent $k$; and
iii) the real continuous variable $s^{k}_{i}\in\mathbb{R}$ denotes arrival time of agent $k$ to task $i$.
The optimization problem is expressed as
@@
Eq. (2) is the scalar objective function that maximizes the convex combination of $R$ objective criteria. Here, $f_{r}$ is $r^{th}$ objective criteria for $r=1,2,...R$. The user-specified parameter $\lambda_{r}$ represents trade-off preference for $r^{th}$ objective criterion. Each objective criterion can depend on task allocation, selected mode, and arrival times.
The constraints of the optimization are categorized into five sets as follows: 1) task assignment constraint, 2) feasibility constraints, 3) energy consumption constraint, 4) time constraints, and 5) domain and preference constraints. These constraints are described next.

### v3

@@ Abstract
[…]
@@ I Introduction and Related Works
Multi-robot teams are widely used in diverse applications such as search and rescue (Liu and Nejat,, 2016; Olson et al.,, 2012), environment mapping (Saeedi et al.,, 2016; Lajoie et al.,, 2022), disaster management response (Phan and Liu,, 2008), marine operations (Thompson and Guihen,, 2019), and agricultural tasks (Das et al.,, 2025). In healthcare settings, these teams can play a vital role in maintaining hygiene to prevent hospital-acquired infections. Disinfection robots offer an effective solution by automating surface disinfection, adding an extra layer of protection. A team of these robots can enhance both spatial and temporal efficiency in maintaining hygiene standards in hospitals (Mehta et al.,, 2023). Coordinating a fleet of disinfection robots requires making two key decisions: i) assigning tasks to robots, and ii) determining the level of disinfection quality for each task. Although higher disinfection quality is preferable, it also requires more time and resources, which may not always be feasible. Task allocation for multi-robot systems has been well-studied in the literature (Korsah et al.,, 2013), but the decision regarding quality of service (QoS) for individual tasks remains largely unexplored. In disinfection applications, QoS is critical, as some surfaces, such as those in patient rooms and ICUs, require higher hygiene standards. This work addresses the challenge of selecting tasks and determining the appropriate disinfection quality/mode for each task.
Multi-Robot Task Allocation (MRTA) algorithms (Gerkey and Matarić,, 2004; Korsah et al.,, 2013) assign tasks to robots to optimize a joint mission objective. A key factor in evaluating these allocations is the QoS of the mission. The target of QoS varies depending on the application. For instance, QoS can refer to the quality of communication between robots and a base station (Khan et al.,, 2022; Wang et al.,, 2023), reduced task completion delays (Botros et al.,, 2023), quantity of area coverage for agricultural survey applications (Barrientos et al.,, 2011), reduced average customer waiting time (Miller and How,, 2017), or optimal worker productivity and reduced fatigue (Calzavara et al.,, 2023). In many cases, QoS competes with other mission objectives, requiring trade-offs among different priorities (Hayat et al.,, 2020; Flushing et al.,, 2017; Calzavara et al.,, 2023; Liang et al.,, 2023). While QoS has been considered to some extent in the MRTA literature, no prior work incorporates QoS as a decision variable specifically for disinfection applications. Introducing QoS as an explicit decision variable enables more precise control over the trade-offs between disinfection quality and other mission objectives. In disinfection scenarios, for example, it allows for optimizing hygiene standards while maximizing task completion within the available mission time.
Agent 1Agent 2Task 1Task 2Task 4Task 3(a)Agent 1Agent 2Task 1Task 2Task 4Task 3(b)Agent 1Agent 2Task 1Task 2Task 4Task 3(c)Mode 1Mode 2Mode 3Missed Task
Figure 1: Traditional Multi-Robot Task Allocation models overlook Quality of Service (QoS) as a decision variable, assuming agents provide a uniform service level, which can lead to inefficiencies in routing and scheduling under time and resource constraints. The visualization compares three cases: (a) Agents operate at the highest service level (Mode 1), ensuring maximum quality but consuming more time and energy, resulting in missed tasks (e.g., Tasks 2 and 3). (b) Agents operate at the lowest service level (Mode 3), completing all tasks but at reduced quality. (c) M3RS incorporates QoS as a decision variable, allowing agents to balance quality and quantity dynamically–choosing from Mode 1, Mode 2 (moderate quality), and Mode 3–ensuring all tasks are addressed at appropriate service levels. Unlike conventional methods, M3RS enables flexible task allocation by integrating QoS into decision-making, making it well-suited for multi-robot applications such as disinfection and cleaning, where balancing quality and quantity is crucial.
Additional decision variables for QoS can further increase the computation complexity of the MRTA problem, which is inherently $\mathcal{NP}$-hard. While heuristics and metaheuristics can provide faster solutions, they often compromise optimality (Braekers et al., 2016b, ). Learning-based approaches require extensive training data and may struggle to generalize (Bogyrbayeva et al.,, 2022). To efficiently handle the added complexity of task-level QoS variables, we propose a clustering-based column generation algorithm. Column generation (Desrosiers and Lübbecke,, 2005) iteratively refines solutions by solving smaller subproblems, allowing for faster convergence to high-quality solutions. The integration of clustering further helps in reducing the search space to find good quality solutions.
To the best of our knowledge, this work presents the first MRTA problem to incorporate task-level QoS using multiple execution modes for quality-dependent multi-robot missions. We introduce the Multi-Robot, Multi-Objective, and Multi-Mode Routing and Scheduling (M3RS) problem, which enables adjustment of task-level QoS to optimize trade-offs among mission objectives. M3RS is particularly relevant for applications where tasks are spatially distributed, have varying service and resource requirements, and involve multiple mission criteria. Each robot can execute tasks in different modes, where each mode dictates task quality, resource consumption, and service time. A visual representation of M3RS compared to the typical MRTA problem is provided in Fig. 1. Traditional MRTA models ignore QoS, while we consider it as a decision variable. We use M3RS for multi-robot disinfection, where both disinfection quality and task throughput are critical. High-quality disinfection, however, increases time and energy costs, limiting the number of tasks performed. M3RS addresses this trade-off by leveraging task-execution modes to balance quality and efficiency. Our contributions are briefly summarized below:
@@
    We propose a novel MRTA problem called Multi-Robot, Multi-Objective, and Multi-Mode Routing and Scheduling (M3RS) that explicitly considers task-level QoS.
@@
[…]
@@
[…]
  4. 4.
    We present simulated and real-world experiments for multi-robot disinfection missions.
@@
II Literature Review: QoS in Task Assignment
QoS has also received attention in the literature on routing and production scheduling. Expósito et al., (2019) considered the timeliness of vehicles as a QoS indicator for vehicle routing. Similarly, QoS has been studied for home healthcare applications, where it involves routing nurses to different patients for healthcare-related services, depending on the patient receiving care from a nurse with the desired experience (Kordi et al.,, 2023; Di Mascolo et al.,, 2021). In supply chain and production planning, QoS has been considered as the existence of minimum stock (Anli et al.,, 2007). In logistics application, customer satisfaction (Liang et al.,, 2023) and optimizing the service times (Gonzalez-R et al.,, 2024) have also been considered as indicators of QoS. Most of these existing works modify the task allocation to achieve different trade-offs between QoS and other objectives. However, there are no works that focus on the deliberate degradation or improvement of QoS offered to an individual task as a means to achieve an overall trade-off for various mission objectives.
Task execution modes refer to distinct operational settings for completing tasks, often chosen based on environmental constraints or desired performance metrics such as tardiness and resource consumption. This concept has been extensively studied in domains like project scheduling and vehicle routing. In project scheduling, execution modes typically represent alternative ways to perform an activity—for instance, allocating more resources to reduce its duration (Hartmann and Briskorn,, 2022; Wkeglarz et al.,, 2011; Tiwari et al.,, 2009). In routing applications, mode selection has been used to choose between transportation methods for healthcare service providers (Braekers et al., 2016a, ), manage hybrid vehicle driving strategies to minimize emissions and costs (Seyfi et al.,, 2022), and optimize battery recharging strategies in electric vehicle fleets (Zhang et al.,, 2024). In MRTA, the use of task execution modes has been limited. Some studies have explored their use to determine whether human-robot collaboration is beneficial for improving time efficiency (Liau and Ryu,, 2022; Ferreira et al.,, 2021; Parreño-Torres et al.,, 2024). However, the use of task execution modes to explicitly control the QoS provided to individual tasks remains largely unexplored in robotic applications. This is particularly important in domains such as multi-robot disinfection, where higher disinfection quality (achieved through higher dosages, for example), has clear physical significance but also incurs greater time and energy costs, potentially reducing overall task throughput. We use task execution modes in M3RS to explore such trade-offs in the context of MRTA.
@@ III Multi-Robot Multi-Objective Multi-Mode Routing and Scheduling
In this section, first, the general optimization model for M3RS problem is presented. Then, we present a column generation (CG) scheme to solve the optimization model with faster computation times. Our novel contribution is the use of clustering heuristics, which helps solve the problem more quickly and find solutions with better objective values. Finally, the formulation is applied to the multi-robot disinfection application.
@@
No Notation Description Comment
1 $\mathcal{A}$ Set of agents. Problem Sets
2 $\mathcal{T}$ Set of tasks.
3 $\mathcal{M}_{i}$ Set of modes for task $i$, where $\mathcal{M}_{i}=\{1,2...M\}$.
4 $R$ Number of objective criteria. Objective Function
5 $f_{r}$ $r^{th}$ objective criterion.
6 $\lambda_{r}$ The scalar preference for $r^{th}$ objective Function.
7 $q_{i\scriptsize{\xrightarrow{}}j}$ Energy consumed for traveling from task $i$ to task $j$. Energy Constraints
8 $q_{\text{idle}}^{k}$ Energy consumption during idling by agent $k$.
9 $q_{i,m}$ Energy consumption of task $i$ in mode $m$.
10 $q_{\text{travel}}$ Average current consumption in traveling at a constant speed.
11 $q_{\text{tc}}$ Average current consumption while completing tasks.
12 $q_{\text{idling}}$ Average current consumption while idling.
13 $Q$ Total energy capacity of an agent.
14 $\sigma_{i,m}$ Time to complete task $i$ in mode $m$. Time Constraints
15 $a_{i},b_{i}]$ Start and end time for task $i$.
16 $T_{H}$ Mission time or time available to complete all tasks.
17 $t_{i,j}$ Travel time from task $i$ to task $j$.
18 $p_{i,m}$ The quality reward of doing task $i$ in mode $m$. Disinfection Problem
19 $p^{max}$ The reward for the highest possible quality of disinfection.
20 $\eta$ Load factor: mission workload relative to the robot’s available capacity.
21 $x^{k}_{i{\xrightarrow{}}j}$ A binary decision variable for if agent $k$ goes from task $i$ to task $j$. Decision Variables
22 $y^{k}_{i,m}$ A binary decision variable for task $i$ done in mode $m$ by agent $k$.
23 $s^{k}_{i}$ A decision variable for arrival time of agent $k$ to task $i$.
@@ III-A Optimization Model
A solution for the M3RS problem determines routes, schedules, and task execution modes for a set of agents $\mathcal{A}$, and a set of tasks $\mathcal{T}$, distributed across a physical space. Refer to Table I for notation and symbols used to formulate the problem. Here, all agents are identical, and only one agent can service each task. We use $i$ and $j$ to denote different tasks in $\mathcal{T}$ and $k$ to denote $k^{th}$ agent in the rest of the notation. The time window for a task $i$ specifies a starting time, $a_{i}$, and finishing time, $b_{i}$, within which the task must be serviced. The entire multi-robot operation must be done in a fixed mission time of $T_{H}$ hours. Further, each task $i$ has an associated mode set $\mathcal{M}_{i}=\{1,2...M\}$, representing available task service modes. Each task can have distinct modes of service available, so each task has a distinct mode set. An agent services the task in one of the modes available in the associated mode set. The selected mode $m\in\mathcal{M}_{i}$ for task $i$ has a corresponding service time ($\sigma_{i,m}$) and energy requirement ($q_{i,m}$) from the agent. Performing a task in a mode with higher service time and resource consumption would lead to lower availability of agents for other tasks. The overall scheduling and allocation mission is optimized for $R$ objective functions.
@@
  1. 1.
    The binary decision variable $x^{k}_{i{\xrightarrow{}}j}\in\{0,1\}$, for assignment and transition of agent $k$ from task $i$ to $j$;
  2. 2.
    The binary decision variable $y^{k}_{i,m}\in\{0,1\}$, for allocation of mode $m$ for task $i$ when executed by agent $k$; and
  3. 3.
    The real continuous variable $s^{k}_{i}\in\mathbb{R}$ denotes arrival time of agent $k$ to task $i$.
With these descriptions, the optimization problem is expressed as follows, optimizing the variables mentioned above:
@@
Eq. (1) is the scalar objective function that maximizes the convex combination of $R$ objective criteria. Here, $f_{r}$ is $r^{th}$ objective criteria for r=1,2,Rr=1,2,...R. The user-specified parameter $\lambda_{r}$ represents the trade-off preference for $r^{th}$ objective criterion. Each objective criterion can depend on task allocation, selected mode, and arrival times.
The constraints of the optimization, described through Eq. (2)-(16), are categorized into five categories, as follows:
  1. 1.
    Task assignment constraint, Eq. (2),
  2. 2.
    Feasibility constraints, Eq. (3)-(7),
  3. 3.
    Energy consumption constraint, Eq. (9),
  4. 4.
    Time constraints, and Eq. (10)-(12), and
  5. 5.
    Domain and preference constraints, Eq. (13)-(16).
These constraints are described next.